OpenStudy (anonymous):

Write a formula to express the following sentence: The sale price is 12% off the original price. Use S for sale price and P for original price to express S as a function of P

OpenStudy (anonymous):

S = P - 0.12P = (1-0.12)P S(P) = 0.88P

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."